OpenStudy (anonymous):

A model rocket is launched vertically with an engine that is ignited at time t = 0. The engine provides an upward acceleration of 30 m/^2 for 2 seconds. Upon reaching its maximum height, the rocket deploys a parachute, and then descends vertically toward the ground. A) Determine the speed of the rocket after the 2s firing of the engine B) What maximum height will the rocket reach? C) At what time after t=0 will the maximum height be reached?

OpenStudy (anonymous):

\[a = 30 + 10 = 40 m/s^2\]

OpenStudy (anonymous):

\[V = u +at\] \[V = 40*2 = 80m/s\]

OpenStudy (anonymous):

\[S = ut + \frac{ 1 }{ 2 }at^2\] \[S = \frac{ 1 }{ 2 } * 40 * 2*2 = 80m\]

OpenStudy (anonymous):

At t=0 the The system is @ rest.....i mean t =0 So.....S = 0m

OpenStudy (anonymous):

@feinne Did u understand..)

OpenStudy (anonymous):